# The NEID Earth Twin Survey. III. Survey Performance After Three Years on Sky

### v1

@@ 3.4 RMS
We measure RMS values of m s-1 in both Run 1 and Run 2 for HD 26965 and HD 38858. As we note in Section 5.4, both stars exhibit moderate RV variations that were originally attributed to planetary companions but have since been determined to be driven by stellar activity. These variations are consistent with the elevated RMS we observe. The RMS value for HD 24496 A rises steeply from just m s-1 to m s-1 during Run 2. There is no coherent rotation signal in the RVs, but this increase is consistent with rising stellar activity levels as traced by the NEID SHK and H indicators, which in turn align with the 5.86-year activity cycle identified by Isaacson et al. (2024) using HIRES data. HD 170657 also appears highly variable, with the RMS just below 3 m s-1 in Run 1 and rising to just above 3 m s-1 in Run 2. No periodic signals are detected in the NEID RVs or activity indicators, but the relatively high SHK,MW value of is consistent with elevated RV variability due to chromospheric activity (Luhn et al., 2020).
@@ 4 RVSearch Analysis of NEID RVs
We use RVSearch (Rosenthal et al., 2021) to perform Bayesian Information Criterion (BIC) model comparisons via an iterative BIC periodogram search of the NEID RVs. The intent of this search is to probe the sensitivity of the NETS data set to known and new periodic signals for each target. In particular, we aim to test whether known planets can be independently recovered with NEID without reference to archival measurements. We therefore limit the analysis to NEID RVs, even though the inclusion of other RV data should in principle lead to a higher recovery rate. Below, we summarize key features of RVSearch that are relevant to our implementation of the algorithm.
The periodogram search is performed over a grid with the same frequency spacing used in Rosenthal et al. (2021), where is the total observing baseline for each target. A BIC periodogram is constructed by stepping through the period grid, fitting a fixed-period sinusoid to the data, and performing a BIC comparison between the and planet models. As in Rosenthal et al. (2021), we adopt a BIC detection threshold corresponding to an empirical false alarm level of 0.1%. At each step in the iterative planet search, if any periodogram peaks exceed this threshold, the most significant of these is used to perform a maximum a posteriori Keplerian fit with all model parameters free. After the solution is determined, an additional planet is added to the model and another iteration of the periodogram search is carried out to compare the and planet models until no more peaks exceed the false-alarm threshold. For each step, the parameters for the maximum likelihood fit of each planet model are left free so that they may converge to an optimal solution as additional planets are identified. The NEID data from Runs 0.5, 1, and 2 are treated as separate time series for the purposes of this analysis, and we allow the model to freely fit the RV zero points for each run.
@@ 5.1 Known Planet Systems - Short Periods
There are four planet hosts in our sample for which all known companions have . These are HD 86728 (1 planet), HD 115617 (2 planets), HD 143761 (4 planets), and HD 168009 (1 planet). For HD 168009, we find that the goodness-of-fit test prefers a linear trend model, while the tests for the remaining stars prefer models without any trend/curvature terms. We find that, depending on the search parameters, either six or seven of these eight known planet signals can be independently detected in the NETS data set. HD 168009 b is detected when a linear trend is included in the fit, but the signal falls just short of the 0.1% false alarm threshold in the BIC periodogram when no trend is included. HD 143761 d is not detected. In addition, we find further evidence to support the Rosenthal et al. (2021) classification of the 121-day RV signal for HD 115617 as a systematic and not a real planet. All signals identified by the search that allows for a trend are listed in Table 3, and signals identified by the search without a trend are listed in Table 4.
@@ 5.1.1 HD 86728
HD 86728 (GJ 376 A) is a slightly evolved G-type star with a wide separation M-dwarf companion, GJ 376 B (Gizis et al., 2000). A 31-day RV signal was first reported by Hirsch et al. (2021) in an analysis of measurements from the HIRES, APF, and the Lick-Hamilton spectrographs. The NEID data presented herein were later used by Gupta et al. (2025) alongside archival APF data to confirm this candidate signal. Gupta et al. (2025) find that HD 86728 b is an exoplanet with orbital period d, projected mass , and an orbital eccentricity consistent with zero.
The signal of HD 86728 b is identified in both of our searches, and the recovered parameters are consistent with the values given by Gupta et al. (2025). No other prominent signals are identified.
@@ 5.1.2 HD 115617
The HD 115617 (61 Vir) system consists of a G-dwarf orbited by two planets first discovered by Vogt et al. (2010). HD 115617 b and c are super-Earth and Neptune-mass planets on nearly circular 4- and 38-day orbits, respectively (Rosenthal et al., 2021). Vogt et al. (2010) reported a third planet, HD 115617 d, at a period of 124 days. Rosenthal et al. (2021) noted that this signal is accompanied by harmonics at days and residual power at 1-year, and following additional analysis of the correlation between the HIRES RVs and PSF parameters, they determined that the signal was a yearly systematic.
Our searches return four signals in both cases. The first two signals are consistent with the orbital parameters of HD 115617 b and c. The other two signals are highly eccentric with periods of 92 and 390 days. The periodograms also show significant peaks at 120 and 180 days. These signals are all likely aliases to a yearly systematic which further corroborates the categorization of HD 115617 d as a false positive by Rosenthal et al. (2021).
@@ 5.1.3 HD 143761
HD 143761 ( CrB), is a near solar mass G-type subgiant with a low metallicity ([Fe/H] = -0.2) and isochrone derived age of Gyrs (Brewer et al., 2023). The system contains four confirmed planets. HD 143761 b is a 40-day Jupiter-mass planet and was one of the first giant planets discovered with RVs (Noyes et al., 1997). HD 143761 c was subsequently discovered by Fulton et al. (2016) with an orbital period of 102 days, and Brewer et al. (2023) added planets d and e which have 282-day and 13-day orbits, respectively.
In both of our searches, three signals were identified and the best-fit orbital parameters are consistent with those of HD 143761 b, c, and e, though we measure a significantly higher eccentricity for HD 143761 e than that reported by Brewer et al. (2023). The searches do not detect the 282-day period of HD 143761 d; the signal is present in the BIC periodograms of the residuals for both searches but it is well below the 0.1% false alarm threshold.
@@ 5.1.4 HD 168009
HD 168009 is a G-dwarf at a distance of 23.3 parsecs hosting a super-Earth discovered via RVs and first reported in Rosenthal et al. (2021). They reported an of 9.53 M⊕, an orbital period of 15.1479 days and eccentricity of 0.121. They also report a 1-year periodic signal that is correlated with instrument systematics and is deemed a false positive.
@@ 5.2 Known Planet Systems - Long Periods
Five of the stars in our sample host at least one planet with . These are HD 95735 (2 planets), HD 154345 (1 planet), HD 186427 (1 planet), HD 190360 (2 planets), and HD 217107 (2 planets). We do not anticipate detecting the five planets with periods that exceed the limit of our search grids but we investigate their impact on the results of performing both blind and informed periodogram searches. The blind searches do not recover any of these planets, as expected. In some cases, the signals exhibit power at a shorter period alias and lead to spurious detections. Still, in both the blind and informed searches, we detect all three of the shorter period planets along with the planet candidate HD 190360 d, which was recently identified by Giovinazzi et al. (2025). Table 3 lists the signals identified by the blind searches, which allows for a trend, and Table 4 lists the results of the informed searches, in which known long-period companion orbits are removed using pre-defined models. Each system and the results of our search are described in detail below.
@@ 5.2.1 HD 95735
HD 95735 (GJ 411) is the lone M-dwarf in the NETS sample. As one of the brightest and nearest M-dwarfs (2.5 pc), HD 95735 has a long history as a target for precision RV surveys. An inner planet, HD 95735 b was first discovered by Díaz et al. (2019) at an orbital period of 13 days, and a longer-period candidate, HD 95735 c was later identified by Rosenthal et al. (2021), though with substantial period uncertainty. A more detailed analysis of the complete set of RVs for this target was carried out by Hurt et al. (2022), who refined the orbital parameters for both planets and confirmed HD 95735 c with an orbital period of 8 yr. Additionally, Hurt et al. (2022) provide constraints on a slightly eccentric candidate signal at 215 days and argue that the planetary case cannot be ruled out despite the classification of this signal as an instrumental systematic by Rosenthal et al. (2021).
In our blind search, a model without any trend/curvature is preferred and three signals are detected. The first has a period of 12.9 days and is consistent with the orbit of HD 95735 b. Eccentric signals at 220 and 654 days are also detected. The first of these is close to the 215-day candidate signal reported by Hurt et al. (2022), but the second has no known comparable signals. Our informed search returns only one signal, which is again the 12.9-day signal corresponding to HD 95735 b. The longer period eccentric signals from the blind search are not identified, suggesting that these were spurious detections caused by the long term trend induced by HD 95735 c.
@@ 5.2.2 HD 154345
HD 154345 is a nearby G-dwarf with a single long-period giant planet companion. Using RV measurements that did not span an entire orbit, Wright et al. (2007) placed initial constraints on a substellar companion, HD 154345 b. The planet was then confirmed by Wright et al. (2008), with a 9-yr orbital period and a minimum mass of close to 1 MJup. The nature of the RV signal was later called into question due to a strong correlation between the RVs and the stellar activity level (Wright, 2016), but several follow up studies have maintained the RV variations to be of planetary origin (Rosenthal et al., 2021; Xiao et al., 2023). Most recently, Giovinazzi et al. (2025) jointly fit archival RVs and Run 1 NETS RVs with astrometric measurements from Hipparcos and Gaia to constrain the 3-D orbital architecture of the system. The astrometry rules out a face-on orbit for HD 154345 b. Giovinazzi et al. (2025) report a dynamical mass constraint of , an orbital period of years, and an eccentricity of .
@@ 5.2.3 HD 186427
HD 186427 (16 Cyg B) is a G-dwarf with both an equal-mass visual binary companion, HD 186408 (16 Cyg A; also a NETS target), and a Jovian planet companion on a long-period (799-day), highly eccentric () orbit (Cochran et al., 1997).
@@ 5.2.4 HD 190360
HD 190360 is a mid-G type star with near-solar mass, temperature, and surface gravity. The system contains two confirmed planets with orbital periods of 2892 days (HD 190360 b; first discovered by Naef et al., 2003) and 17 days (HD 190360 c; first discovered by Vogt et al., 2005). HD 190360 b is a Jupiter analog, with a true dynamical mass measurement of based on a combination of astrometric and RV measurements, including Run 1 NEID data (Giovinazzi et al., 2025). The astrometric signal of HD 190360 c is below the precision and baseline of the astrometric acceleration measurements, leading to a poor constraint on the mutual inclination of the system. Giovinazzi et al. (2025) also report a candidate signal in the NEID RV residuals with a period of 90 days; this signal had previously been identified but was attributed to a systematic alias by Rosenthal et al. (2021).
@@ 5.2.5 HD 217107
HD 217107 is a G-dwarf hosting two planets, HD 217107 b and c, originally discovered by Fischer et al. (1999) and Vogt et al. (2005), respectively. As with the HD 190360 system, the outer Jupiter analog (HD 217107 c; days) has a dynamical mass measurement based on a combined RV and astrometric data set (Giovinazzi et al., 2025), but the orbit of the inner planet (HD 217107 b; days) does not induce a measurable astrometric signal.
@@ 5.3.1 HD 10700
HD 10700, better known as Ceti, has a contentious history of radial velocity detected planet candidates. Most recently, Feng et al. (2017) reported four planet candidates with sub-m s-1 semi-amplitudes at orbital periods of 20 ( m s-1), 49 ( m s-1), 163 ( m s-1), and 636 ( m s-1) days. The goodness-of-fit test prefers a linear trend of m s-1 yr-1, and in both search cases we detect no significant signals at these periods. However, both searches do identify a signal at 123 days with a semi-amplitude of 1.3 m s-1 and eccentricity . We do not see a peak at this period in the GLS periodogram of any of our activity indicators.
@@ 5.3.3 HD 4628
HD 4628 is an early K-dwarf with a 38.5-day rotation period (Brandenburg et al., 2017). The goodness-of-fit test prefers a linear trend term of m s-1 yr-1. In this case, a 14.7-day signal is identified with a semi-amplitude of 1.14 m/s and an eccentricity of 0.38. In the fit that does not include a trend/curvature term, the strongest peak in the periodogram is at 14.8 days but it does not exceed the false-alarm threshold and is not formally detected. There are no corresponding periodogram peaks for any of the NEID activity indicators, but we do observe a linear correlation between the RV and SHK measurements which can be attributed to rotational modulation and the long-term stellar activity cycle.

### v2

@@ 3.4 RMS
We measure RMS values of m s-1 in both Run 1 and Run 2 for HD 26965 and HD 38858. As we note in Section 5.4, both stars exhibit moderate RV variations that were originally attributed to planetary companions but have since been determined to be driven by stellar activity. These variations are consistent with the elevated RMS we observe. The RMS value for HD 24496 A rises steeply from just m s-1 to m s-1 during Run 2. There is no coherent rotation signal in the RVs, but this increase is consistent with rising stellar activity levels as traced by the NEID SHK and H indicators, which in turn align with the 5.86-year activity cycle identified by H. Isaacson et al. (2024) using HIRES data. HD 170657 also appears highly variable, with the RMS just below 3 m s-1 in Run 1 and rising to just above 3 m s-1 in Run 2. No periodic signals are detected in the NEID RVs or activity indicators, but the relatively high SHK,MW value of is consistent with elevated RV variability due to chromospheric activity (J. K. Luhn et al., 2020).
@@ 4 RVSearch Analysis of NEID RVs
We use RVSearch (L. J. Rosenthal et al., 2021) to perform Bayesian Information Criterion (BIC) model comparisons via an iterative BIC periodogram search of the NEID RVs. The intent of this search is to probe the sensitivity of the NETS data set to known and new periodic signals for each target. In particular, we aim to test whether known planets can be independently recovered with NEID without reference to archival measurements. We therefore limit the analysis to NEID RVs, even though the inclusion of other RV data should in principle lead to a higher recovery rate. Below, we summarize key features of RVSearch that are relevant to our implementation of the algorithm.
The periodogram search is performed over a grid with the same frequency spacing used in L. J. Rosenthal et al. (2021), where is the total observing baseline for each target. A BIC periodogram is constructed by stepping through the period grid, fitting a fixed-period sinusoid to the data, and performing a BIC comparison between the and planet models. As in L. J. Rosenthal et al. (2021), we adopt a BIC detection threshold corresponding to an empirical false alarm level of 0.1%. At each step in the iterative planet search, if any periodogram peaks exceed this threshold, the most significant of these is used to perform a maximum a posteriori Keplerian fit with all model parameters free. After the solution is determined, an additional planet is added to the model and another iteration of the periodogram search is carried out to compare the and planet models until no more peaks exceed the false-alarm threshold. For each step, the parameters for the maximum likelihood fit of each planet model are left free so that they may converge to an optimal solution as additional planets are identified. The NEID data from Runs 0.5, 1, and 2 are treated as separate time series for the purposes of this analysis, and we allow the model to freely fit the RV zero points for each run.
@@ 5.1 Known Planet Systems - Short Periods
There are four planet hosts in our sample for which all known companions have . These are HD 86728 (1 planet), HD 115617 (3 planets), HD 143761 (4 planets), and HD 168009 (1 planet). For HD 168009, we find that the goodness-of-fit test prefers a linear trend model, while the tests for the remaining stars prefer models without any trend/curvature terms. We find that, depending on the search parameters, either six or seven of these nine known planet signals can be independently detected in the NETS data set. HD 168009 b is detected when a linear trend is included in the fit, but the signal falls just short of the 0.1% false alarm threshold in the BIC periodogram when no trend is included. HD 143761 d and HD 115617 d are not detected, though there is some disagreement in the literature as to whether this latter planet is real (L. J. Rosenthal et al., 2021; M. Cretignier et al., 2023). All signals identified by the search that allows for a trend are listed in Table 3, and signals identified by the search without a trend are listed in Table 4.
@@ 5.1.1 HD 86728
HD 86728 (GJ 376 A) is a slightly evolved G-type star with a wide separation M-dwarf companion, GJ 376 B (J. E. Gizis et al., 2000). A 31-day RV signal was first reported by L. A. Hirsch et al. (2021) in an analysis of measurements from the HIRES, APF, and the Lick-Hamilton spectrographs. The NEID data presented herein were later used by A. F. Gupta et al. (2025) alongside archival APF data to confirm this candidate signal. A. F. Gupta et al. (2025) find that HD 86728 b is an exoplanet with orbital period d, projected mass , and an orbital eccentricity consistent with zero.
The signal of HD 86728 b is identified in both of our searches, and the recovered parameters are consistent with the values given by A. F. Gupta et al. (2025). No other prominent signals are identified.
@@ 5.1.2 HD 115617
The HD 115617 (61 Vir) system consists of a G-dwarf orbited by 3 planets first discovered by S. S. Vogt et al. (2010). HD 115617 b and c are super-Earth and Neptune-mass planets on nearly circular 4- and 38-day orbits, respectively (L. J. Rosenthal et al., 2021). S. S. Vogt et al. (2010) reported a third planet, HD 115617 d, at a period of 124 days, which was also independently confirmed by K. Laliotis et al. (2023) and M. Cretignier et al. (2023). However, L. J. Rosenthal et al. (2021) noted that this signal is accompanied by harmonics at days and residual power at 1-year, and following additional analysis of the correlation between the HIRES RVs and PSF parameters, they determined that the signal was a yearly systematic.
Our searches return four signals in both cases. The first two signals are consistent with the orbital parameters of HD 115617 b and c. The other two signals are highly eccentric with periods of 92 and 390 days. The periodograms also show significant peaks at 120 and 180 days. These signals are all likely aliases of the 124-day orbit of HD 115617 d, but it is not clear whether this is a false positive, as suggested by L. J. Rosenthal et al. (2021), or a real planet signal.
@@ 5.1.3 HD 143761
HD 143761 ( CrB), is a near solar mass G-type subgiant with a low metallicity ([Fe/H] = -0.2) and isochrone derived age of Gyrs (J. M. Brewer et al., 2023). The system contains four confirmed planets. HD 143761 b is a 40-day Jupiter-mass planet and was one of the first giant planets discovered with RVs (R. W. Noyes et al., 1997). HD 143761 c was subsequently discovered by B. J. Fulton et al. (2016) with an orbital period of 102 days, and J. M. Brewer et al. (2023) added planets d and e which have 282-day and 13-day orbits, respectively.
In both of our searches, three signals were identified and the best-fit orbital parameters are consistent with those of HD 143761 b, c, and e, though we measure a significantly higher eccentricity for HD 143761 e than that reported by J. M. Brewer et al. (2023). The searches do not detect the 282-day period of HD 143761 d; the signal is present in the BIC periodograms of the residuals for both searches but it is well below the 0.1% false alarm threshold.
@@ 5.1.4 HD 168009
HD 168009 is a G-dwarf at a distance of 23.3 parsecs hosting a super-Earth discovered via RVs and first reported in L. J. Rosenthal et al. (2021). They reported minimum mass of M⊕, an orbital period of 15.1479 days and eccentricity of 0.121. They also report a 1-year periodic signal that is correlated with instrument systematics and is deemed a false positive.
@@ 5.2 Known Planet Systems - Long Periods
Five of the stars in our sample host at least one planet with . These are HD 95735 (2 planets), HD 154345 (1 planet), HD 186427 (1 planet), HD 190360 (2 planets), and HD 217107 (2 planets). We do not anticipate detecting the five planets with periods that exceed the limit of our search grids but we investigate their impact on the results of performing both blind and informed periodogram searches. The blind searches do not recover any of these planets, as expected. In some cases, the signals exhibit power at a shorter period alias and lead to spurious detections. Still, in both the blind and informed searches, we detect all three of the shorter period planets along with the planet candidate HD 190360 d, which was recently identified by M. R. Giovinazzi et al. (2025). Table 3 lists the signals identified by the blind searches, which allows for a trend, and Table 4 lists the results of the informed searches, in which known long-period companion orbits are removed using pre-defined models. Each system and the results of our search are described in detail below.
@@ 5.2.1 HD 95735
HD 95735 (GJ 411) is the lone M-dwarf in the NETS sample. As one of the brightest and nearest M-dwarfs (2.5 pc), HD 95735 has a long history as a target for precision RV surveys. An inner planet, HD 95735 b was first discovered by R. F. D´ıaz et al. (2019) at an orbital period of 13 days, and a longer-period candidate, HD 95735 c was later identified by L. J. Rosenthal et al. (2021), though with substantial period uncertainty. A more detailed analysis of the complete set of RVs for this target was carried out by S. A. Hurt et al. (2022), who refined the orbital parameters for both planets and confirmed HD 95735 c with an orbital period of 8 yr. Additionally, S. A. Hurt et al. (2022) provide constraints on a slightly eccentric candidate signal at 215 days and argue that the planetary case cannot be ruled out despite the classification of this signal as an instrumental systematic by L. J. Rosenthal et al. (2021).
In our blind search, a model without any trend/curvature is preferred and three signals are detected. The first has a period of 12.9 days and is consistent with the orbit of HD 95735 b. Eccentric signals at 220 and 654 days are also detected. The first of these is close to the 215-day candidate signal reported by S. A. Hurt et al. (2022), but the second has no known comparable signals. Our informed search returns only one signal, which is again the 12.9-day signal corresponding to HD 95735 b. The longer period eccentric signals from the blind search are not identified, suggesting that these were spurious detections caused by the long term trend induced by HD 95735 c.
@@ 5.2.2 HD 154345
HD 154345 is a nearby G-dwarf with a single long-period giant planet companion. Using RV measurements that did not span an entire orbit, J. T. Wright et al. (2007) placed initial constraints on a substellar companion, HD 154345 b. The planet was then confirmed by J. T. Wright et al. (2008), with a 9-yr orbital period and a minimum mass of close to 1 MJup. The nature of the RV signal was later called into question due to a strong correlation between the RVs and the stellar activity level (J. T. Wright, 2016), but several follow up studies have maintained the RV variations to be of planetary origin (L. J. Rosenthal et al., 2021; G.-Y. Xiao et al., 2023). Most recently, M. R. Giovinazzi et al. (2025) jointly fit archival RVs and Run 1 NETS RVs with astrometric measurements from Hipparcos and Gaia to constrain the 3-D orbital architecture of the system. The astrometry rules out a face-on orbit for HD 154345 b. M. R. Giovinazzi et al. (2025) report a dynamical mass constraint of , an orbital period of years, and an eccentricity of .
@@ 5.2.3 HD 186427
HD 186427 (16 Cyg B) is a G-dwarf with both an equal-mass visual binary companion, HD 186408 (16 Cyg A; also a NETS target), and a Jovian planet companion on a long-period (799-day), highly eccentric () orbit (W. D. Cochran et al., 1997).
@@ 5.2.4 HD 190360
HD 190360 is a mid-G type star with near-solar mass, temperature, and surface gravity. The system contains two confirmed planets with orbital periods of 2892 days (HD 190360 b; first discovered by D. Naef et al., 2003) and 17 days (HD 190360 c; first discovered by S. S. Vogt et al., 2005). HD 190360 b is a Jupiter analog, with a true dynamical mass measurement of based on a combination of astrometric and RV measurements, including Run 1 NEID data (M. R. Giovinazzi et al., 2025). The astrometric signal of HD 190360 c is below the precision and baseline of the astrometric acceleration measurements, leading to a poor constraint on the mutual inclination of the system. M. R. Giovinazzi et al. (2025) also report a candidate signal in the NEID RV residuals with a period of 90 days; this signal had previously been identified but was attributed to a systematic alias by L. J. Rosenthal et al. (2021).
@@ 5.2.5 HD 217107
HD 217107 is a G-dwarf hosting two planets, HD 217107 b and c, originally discovered by D. A. Fischer et al. (1999) and S. S. Vogt et al. (2005), respectively. As with the HD 190360 system, the outer Jupiter analog (HD 217107 c; days) has a dynamical mass measurement based on a combined RV and astrometric data set (M. R. Giovinazzi et al., 2025), but the orbit of the inner planet (HD 217107 b; days) does not induce a measurable astrometric signal.
@@ 5.3.1 HD 10700
HD 10700, better known as Ceti, has a contentious history of RV-detected planet candidates. Most recently, F. Feng et al. (2017) reported four planet candidates with sub-m s-1 semi-amplitudes at orbital periods of 20 ( m s-1), 49 ( m s-1), 163 ( m s-1), and 636 ( m s-1) days. The goodness-of-fit test prefers a linear trend of m s-1 yr-1, and in both search cases we detect no significant signals at these periods. However, both searches do identify a signal at 123 days with a semi-amplitude of 1.3 m s-1 and eccentricity . Ceti was recently shown to have a low stellar inclination, such that we are observing the star nearly pole-on relative to the rotation axis (M. Korolik et al., 2023). At this viewing angle it is unlikely for rotational modulation to affect the RVs at the m s-1 level, but the 123-day signal could be an alias of the -day rotation period as spots move in and out of view on the stellar limb. We do not see a peak at either of these periods in the GLS periodogram of any of our activity indicators.
@@ 5.3.3 HD 4628
HD 4628 is an early K-dwarf with a 38.5-day rotation period (A. Brandenburg et al., 2017). The goodness-of-fit test prefers a linear trend term of m s-1 yr-1. In this case, a 14.7-day signal is identified with a semi-amplitude of 1.14 m/s and an eccentricity of 0.38. In the fit that does not include a trend/curvature term, the strongest peak in the periodogram is at 14.8 days but it does not exceed the false-alarm threshold and is not formally detected. There are no corresponding periodogram peaks for any of the NEID activity indicators, but we do observe a linear correlation between the RV and SHK measurements which can be attributed to rotational modulation and the long-term stellar activity cycle.